OpenStudy (anonymous):

Solve. 2x2 – 3x + 6 = 0

OpenStudy (anonymous):

4-3x+6=0 4+6-3x=0 10-3x=0 transfer 3x to the other side so: 10=3x divide both sides by 3 so: 10/3 = x

hero (hero):

hbaldon, that isn't even close to right

OpenStudy (anonymous):

wait.. is that 2x^2?

hero (hero):

Yes

OpenStudy (anonymous):

oh. sorry. ok ok. so that's quadratic. use quadratic formula :D

OpenStudy (anonymous):

i thought it's 2 multiplied to 2. sorry, my bad. :)

OpenStudy (anonymous):